OpenStudy (anonymous):

Find the exact value of the logarithm without using a calculator. log2(8) + log2(16)

OpenStudy (anonymous):

i know the answer, i want to know hwo to get to it

OpenStudy (anonymous):

ansewr is 5

OpenStudy (anonymous):

are these to the base 2 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use the properties of log, log(a) +log(b)=log(ab) so using it here we get log(8)+log(16)=log(8*16) = log(\[2^{7}\]=7 loga(a) =1.

OpenStudy (anonymous):

the answer is 7

OpenStudy (anonymous):

no, the answer is 5.

OpenStudy (anonymous):

oh my bad, different problem

OpenStudy (anonymous):

\[log_2(8) + log_2(16) = log_2(2^3) + log_2(2^4) = 3 + 4 = 7\]

OpenStudy (anonymous):

polpak keep doing it please

OpenStudy (anonymous):

i like the way you doing it

OpenStudy (anonymous):

Well brackett's method works better when you don't have exact powers of the base.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but how you can figure out that 128=2^7?

OpenStudy (anonymous):

Yeah, that's tough unless you know a lot of powers of 2.

myininaya (myininaya):

\[\ln 128=\ln 2^x\] \[\ln 128=x\ln2\] \[\frac{\ln128}{\ln2}=x\] \[x=\frac{\ln 2^7}{\ln2}=\frac{7*\ln2}{\ln2}=7\]